Status Report

NASA Space Station On-Orbit Status 4 Aug 2004

By SpaceRef Editor
August 4, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 4 Aug 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below. 

After yesterday’s successful EVA-10 and a sleep period that ended this morning at 2:00am EDT, the crew worked post-spacewalk tasks on their Orlan-M spacesuits, checking the water quantity in the cooling water bladders, with refills as required, and removing the BK-3 oxygen (O2) tanks from the backpacks and the batteries from the BRTA radio telemetry units. 

Gennady collected the radiation measurements taken by the “Pille-MKS” dosimeter sensors in the two Orlan suits and the static background monitor in the Service Module (SM).  The three sensors were then deployed at their regular positions in the Russian segment (RS).   [“Pille” has ten sensors normally situated at various locations in the RS (port cabin window, stbd cabin window, ASU toilet facility, control panel, etc.).  Dosage values are called down or downlinked via Regul Paket/Email or OCA.]

The CDR set up the Orlan battery-charging unit (ZU-S), then initiated discharge of the two 825M3 Orlan backpack batteries for storage.

Padalka set up each of the spacesuits for drying, in turn.  Afterwards, the Orlans and their BSS interface control units were returned to their storage in the DC1 docking compartment.

Gennady also reactivated the new “Sputnik-SM” ham radio hardware in the SM, then stowed the Urolux hardware used by the crew last night for the obligatory post-EVA session of the Russian crew health-monitoring program’s MO-9/Biochemical Urinalysis.

FE/SO Mike Fincke had an additional 15 min. for transferring the U.S. EVA tools back to the U.S. segment (USOS) for stowage.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

At 6:30am, the crew debriefed with ground specialists on the spacewalk via S-band.  For the debriefing, MCC-H had uplinked a write-up on the temporary loss of attitude during the EVA.   [The possibility that the U.S. CMGs (control moment gyros) may become saturated by an as-yet not fully explained disturbance torque on the ISS during the time the DC-1 hatch was open had been considered and prepared for on the basis of previous data from EVA-9 and EVA-9A last month.  Thus, the occurrence of LOAC (loss of attitude control) as a result of the saturation during the period the SM thrusters were inhibited was anticipated.  With the solar arrays producing less power, and because CMG saturation occurred during the second half of the EVA (i.e., with most of the tasks done), MCC-H commanded a load shed (controlled reduction of power usage) while the crew continued the EVA in free drift.  However, after the load-shed initiation ISS unexpectedly lost S-band communications.  At that point, Houston and Moscow agreed that during the next Russian ground site pass, Moscow would ask Padalka & Fincke to translate away from SM aft so that attitude control could be regained using SM thrusters in case the loss of S-Band was caused by the station drifting into a bad attitude for comm.  In parallel, Houston sent commands via Russian assets to activate S-Band String 1, which was in a “warm back up” mode.  S-band comm was immediately recovered, attitude control was handed back to the CMGs, and the EVA was completed successfully.]

FE/SO Mike Fincke activated the EXPRESS rack 3 (ER3) laptop, then loaded it with a software upgrade for the ER3 ICA (interface controller unit).  The upgrade ended at ~7:15am, and ER3 was turned off again later in the day after a data downlink by POIC (Payload Operations & Integration Center).

Fincke demated and removed the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab and Cupola RWSs (robotics work stations) that were used to support video camera coverage of yesterday’s EVA.

The FE also performed the post-EVA reboot of all operational PCS (portable computer system) laptops and restarted the OCA comm router laptop.

Mike conducted the (almost-)regular weekly task of filling out the FFQ (food frequency questionnaire), which keeps a log of his nutritional intake over time on special MEC (medical equipment computer) software.

Padalka completed the daily routine inspection of the SM’s SOZh life support system, including the periodic inspection of the BRPK air/liquid condensate separator apparatus as well as the periodic checkout of the IP-1 airflow sensors in the various RS hatchways and FGB-to-Node tunnel (last time done: 7/21), while Fincke conducted the regular routine status checkup of the autonomous ISS-7 Lab payloads (PCG-STES010, MAMS).

Mike and Gennady also performed their full regimen of physical exercise on VELO with force loader, RED (resistive exercise device) and TVIS (treadmill with vibration isolation and stabilization).

Working off the Russian task list, the CDR will be conducting another run of the Russian Uragan earth-imaging program, using the Kodak 760 DSC (digital still camera) with 800mm-lens from SM windows #9.   [Today’s observation features imagery of the Lake Baikal coastal area and Amur river flood plain on both sides of track (Triple crossing of Amur river), Sakhalin island coastal area, Crimean mountains, the Gulf of Taganrog, the Eastern Volga shore, the Southern Urals, Altai glaciers, coastline of France from Marseille and Toulon to Nice and Monaco, the cities of Vienna and Bratislava, etc.]

The crew was given viewing data for video recording of three possible launches of Equis II sounding rockets from Kwajalein Atoll.   [The first rocket is an instrument carrier that is launched to the West. The second rocket, also launched to the West along the same azimuth as the instrument carrier, is designed to release Trimethyl Aluminum (TMA) during portions of both the up-leg and down-leg.  The third rocket, also for TMA release, will be launched to the North.]

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

During the day, TsUP/Moscow is conducting tests of the automated passive Kurs-P rendezvous and approach radar system on the SM & FGB as well as on the Soyuz external video camera.

The station continues to fly in XPOP attitude (X-axis perpendicular to orbit plane), pitch: 0.8 deg, yaw: -8.0 deg, roll: 0 deg.

Major upcoming events:

  • Progress 15P launch — 8/11 (1:01am EDT);
  • Progress 15P docking — 8/14 (2:05am EDT).

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 1:29am EDT [= epoch]):

  • Mean altitude — 357.4 km
  • Apogee height — 361.6 km
  • Perigee height — 353.2km
  • Period — 91.7 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0006187
  • Solar Beta Angle — -25.3 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32586

SpaceRef staff editor.